MKS Inc., produces meter sticks that have a target length of 100 centimeters with upper and lower

specification limits of 100.05 and 99.95 centimeters respectively. Their existing process produces

meter sticks with an average length of 100.00 centimeters and a standard deviation of 0.015

centimeters.

The value of the process capability index to two decimal places is ?
Answer #1

Upper specification limit (USL) = 100.05

Lower specification limit (LSL) = 99.95

Mean = 100.00

Standard deviation (SD) = 0.015

Cpl = (Mean - LSL) / (3 x SD) = (100.00-99.95)/(3x0.015) = 0.05/0.045 = 1.11

Cpu = (USL - Mean) / (3 x SD) = (100.05-100.00)/(3x0.015) = 0.05/0.045 = 1.11

Process capability index (Cpk) = Minimum of (Cpu, Cpl)

= Minimum of (1.11 , 1.11)

= 1.11